OpenStudy (anonymous):

how do l solve 16x^2=49

OpenStudy (anonymous):

divide by 16, then take the square root don't forget the \(\pm\)

OpenStudy (anonymous):

Bring the 49 over to the left. 16x^2 -49 =0 Recall that \[a ^{2} - b ^{2} = (a+b)(a-b)\] so what is the 4^2 and 7^2 = ?